# Bringing Generative AI to Adaptive Learning in Education

### v2

@@ Abstract
The recent surge in generative AI technologies, such as large language models and diffusion models, have boosted the development of AI applications in various domains, including science, finance, and education. Concurrently, adaptive learning, a concept that has gained substantial interest in the educational sphere, has proven its efficacy in enhancing students’ learning efficiency. In this position paper, we aim to shed light on the intersectional studies of these two methods, which combine generative AI with adaptive learning concepts. By presenting discussions about the benefits, challenges, and potentials in this field, we argue that this union will contribute significantly to the development of the next-stage learning format in education.
@@ 1 Introduction
Generative AI (GenAI) refers to artificial intelligence models that generate new content, including text, images, audio, and video, based on the patterns and information they have learned from the given training data (Cao et al., 2023). Unlike the traditional Machine Learning (ML) algorithms, which focus on analyzing and interpreting data, GenAI is designed to create new, original outputs and thus able to solve more challenging problems. With the recent success of GenAI technologies, such as large language models and diffusion models, in generating human-like outputs to a wide range of problems in Natural Language Processing (NLP) (Bubeck et al., 2023) and Computer Vision (CV) (Ramesh et al., 2022), the development of applying GenAI to real-world problems is rapidly increasing, which spreads from science (Walters & Murcko, 2020; Lopez et al., 2020), finance (Rane, 2023; Brühl, 2023) to education (Cooper, 2023; Baidoo-Anu & Ansah, 2023).
@@
Why This Position Paper? The overwhelming performance of GenAI over ML algorithms in challenging data analysis problems provides new opportunities to current ML-based AL systems. By leveraging the advanced capabilities of GenAI in creating human-level responses to problems like reasoning, the new AL systems will be able to finish more complicated AL tasks, such as dynamic learning path planning. Although some pioneering works have started to explore the usage of GenAI for education purposes (Dan et al., 2023; Xiao et al., 2024), there is no systematic review and discussion focusing on bringing GenAI to AL problems. As adaptive learning is increasingly recognized as a significant and promising direction for the future of education (Park & Lee, 2013), in this position paper, we propose to review the related studies and call attention to the opportunities in combining GenAI with AL. Furthermore, by conducting comprehensive discussions on the benefits, challenges, and threats, we hope to shed light on future research in this direction and push adaptive learning into the next chapter.
Our Contributions.
@@
Our position in this paper is that bringing GenAI to AL will present both advantages and challenges to current ML-based AL systems, creating numerous opportunities and original topics for future educational development. To support this claim, we (1) conduct a comprehensive literature review on key ML contributions to adaptive learning, (2) explore the potential benefits and summarize ongoing industrial practices, (3) present in-depth discussions about the challenges and opportunities from the pedagogical perspective, and further extend these considerations to their broader educational impacts.
Related Works.
@@
The impressive performance of GenAI, especially the Large Language Models (LLMs), in generating human-like responses to complicated requests (Bubeck et al., 2023) encourages the recent research in exploring GenAI for education purpose (Dan et al., 2023; Kieser et al., 2023). To provide researchers with a broad overview of the domain, numerous exploratory and survey papers have been proposed. For example, Qadir (2023); Rahman & Watanobe (2023) and Rahman & Watanobe (2023) conclude the applications of ChatGPT to engineering education by analyzing the responses of ChatGPT to the related pedagogical questions. Jeon & Lee (2023) and Mogavi et al. (2023) collect the opinions from different ChatGPT user groups, e.g., educator, learner, researcher, through in-person interviews, online post replies and user logs, and conclude the practical applications of LLMs in education scenarios. Baidoo-Anu & Ansah (2023) and Zhang & Tur (2023) focus on the literature review over the published papers and summarize the progress of the area with structured tables. Although the above works have covered a wide range of existing applications of GenAI in education scenario and provided their long-term visions for future studies, we argue that our position paper is still valuable as none of them focuses on the direction of bringing GenAI to AL. Compared to the general problems such as applying GenAI for education, we propose to present a more concentrated summarization and in-depth discussion toward the adaptive learning topic in this paper.
@@ 2.2 Generative AI (GenAI)
The origins of Generative AI can be traced back to the mid-1950s with the emergence of the concepts of artificial intelligence and machine learning. Unlike discriminative modeling, which has been the driving force of most progress in artificial intelligence, generative modeling problems are generally more difficult to tackle. Contributing to the rise of deep learning (LeCun et al., 2015), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Variational Auto-encoders(VAEs) (Kingma & Welling, 2013) first revolutionized image generation and became a cornerstone in the field of GenAI. Moreover, the recent emergence of Transformer (Vaswani et al., 2017), has marked another significant evolution in GenAI within the field of NLP, which has propelled GenAI into a new era. LLMs, such as OpenAI’s GPT series, are the most representative ones of these GenAI models. By taking advantage of the vast number of parameters, extensive pre-training text corpus, and advanced fine-tuning methods (Ouyang et al., 2022), LLMs have shown impressive capabilities in understanding context, generating coherent and contextually relevant responses (Bubeck et al., 2023). Apart from that, further studies (Brown et al., 2020) discover the strong few-shot (in-context) learning ability of LLMs, and such findings encourage the revolution in NLP solutions from task-specific methods to foundation models (Zhou et al., 2023a). Despite the above progress, Generative Diffusion Models (GDMs) (Ho et al., 2020; Song et al., 2020) also bring improvements to the generation quality of images (Dhariwal & Nichol, 2021) and audio signals (Kong et al., 2020). Finally, by connecting GenAI models from different modalities, multi-modal GenAI models are proposed and have achieved promising results in directions like text-to-image (Ramesh et al., 2022), text-to-audio (Popov et al., 2021) and text-to-video generations (Wang et al., 2023a).
3 Machine Learning in Adaptive Learning
@@
Machine learning plays a crucial role in adaptive learning by enabling the system to analyze learners’ behavior, build learning material hierarchies, and provide personalized adaptions to learners. In the following, we summarize three pivotal ML algorithms used by AL systems.
3.1 Knowledge Tracing (KT)
Knowledge proficiency profile is the basis for AL systems and its diagnosing process is well-known as the Knowledge Tracing (KT) problem in literature (Anderson et al., 1990). Generally speaking, the goal of KT is to track the learning progress of students, and use it to instruct the following adjustments toward the students’ learning contents (Abdelrahman et al., 2023). However, as proficiency is not an observable variable, KT methods commonly reformulate the problem as predicting student’s next response in the diagnosis test (Villano, 1992). Specifically, given a student’s past exercising interactions $X$ in a system, the goal of KT is to model student exercise interaction procedure to predict the student’s future performance at timestamp $t+1$, i.e., predicting the probability $p(r_{r+1}=1|e_{t+1},X)$, where $r_{t+1}\in\{0,1\}$ indicates the student’s answer correctness on new exercise $e_{t+1}$. The diagram of knowledge tracing is shown in Fig. 2.
Figure 2: Knowledge Tracing Diagram.
3.1.1 Traditional KT
To achieve that goal, traditional researches developed two series of approaches for KT: Bayesian Knowledge Tracing (BKT) and Factor Analysis Models (FAM). BKT was motivated by the concepts of mastery learning (Corbett, 2000) and commonly uses a probabilistic graphical model, such as Hidden Markov Model (Baker et al., 2008) and Bayesian Belief Network (Villano, 1992) to trace the students’ changing knowledge states as they practice skills. Corbett & Anderson (1994) introduce the first BKT model, which models student’s knowledge state with a binary variable $\{unlearned,learned\}$ and considers transitions from the unlearned state to the learned state. Note that a student may make a mistake while in a learned state or guess correctly in an unlearned state, BKT includes the slip probability $p(S)$ and guess probability $p(G)$ to capture that. In particular, the standard BKT can be expressed as:
$p(r_{n}=1)=p(L_{n})*(1-p(S))+(1-p(L_{n}))*p(G),$
where $r_{n}$ is student’s answer correctness on the $n$-th question and $p(L_{n})$ is the posterior estimation of student’s proficiency, which be estimated by Bayesian formula iteratively:
$p(L_{n})=p(L_{n-1})+(1-p(L_{n-1}))*p(T)$
$p(L_{n-1})=\begin{cases}\frac{p(L_{n-1})*(1-p(S)}{p(L_{n-1})*(1-p(S))+(1-p(L_{n-1}))*p(G)},&r_{n}=1\\ \frac{p(L_{n-1})*p(S)}{p(L_{n-1})*p(S)+(1-p(L_{n-1}))*(1-P(G))},&r_{n}=0\end{cases}$
On the other hand, FAM is theoretically supported by the Item Response Theory (IRT) (Harvey & Hammer, 1999), which estimates student performance by learning a function, usually a logistic function, based on various factors in a population of students who solve a set of problems. The probability $p_{ij}$ that a student $i$ correctly answers a question $j$ is defined as:
$p_{ij}=\mathcal{L}(\theta_{i}-b_{j})=\frac{\exp^{(\theta_{i}-b_{j})}}{1+\exp^{(\theta_{i}-b_{j})}},$
where $b_{j}$ is the difficulty of question $j$ and $\theta_{i}$ is the student $i$’s ability. To estimate $\theta_{i}$, FAM methods take different factors into their consideration. For example, the Additive Factor Model (AFM) (Cen et al., 2008) counts the frequency that a student meets with different skills and uses a linear combination over them to estimate $\theta$. Pavlik Jr et al. (2009) and Thai-Nghe et al. (2012) further add student’s correctness records on the past questions and second order interactions between different skills to increase the complexity of the FAM model for achieving better KT results.
3.1.2 Deep Learning based KT
Inspired by the success of Deep Learning (DL) (LeCun et al., 2015), recent researches on KT have widely adopted deep learning techniques. Deep Knowledge Tracing (DKT) (Piech et al., 2015) was the pioneering study that models KT as a sequence prediction problem and uses Recurrent Neural Network (RNN) (Lipton et al., 2015) to predicate the probability of correctly answering a question at each time step $t$. Furthermore, by taking advantage of the strong fitting capabilities of deep learning methods, recent deep learning based KT models propose a variety of network designs to take more factors into account. For example, memory-augmented KT traces complex Knowledge Concepts (KCs) learned by students and extended DKT by augmenting an external memory structure (Zhang et al., 2017; Abdelrahman & Wang, 2019), attentive KT attempts to incorporate an attention mechanism to mitigates prior model’s limitation in treating all questions in a sequence of interactions equally important (Ghosh et al., 2020; Pandey & Karypis, 2019), graph-based KT leverages graph representation learning techniques (Nakagawa et al., 2019; Tong et al., 2020) to capture such relational structures between questions and their KCs, text-Aware KT leverages the textual features of questions to enhance the performance (Su et al., 2018; Liu et al., 2019). In general, deep learning based KT models have demonstrated their great potential in solving the KT problem.
3.2 Knowledge Concepts Structure Construction
The hierarchical structure of knowledge concepts and learning materials helps AL systems in identifying the appropriate learning contents for students in an efficient manner (Adjei et al., 2016). Because of the heavy manual workload in organizing the huge amount of learning content into a complete knowledge concept framework, following ML methods are widely adopted by AL systems and the diagram is shown as Fig. 3.
Figure 3: Knowledge Concepts Structure Construct Diagram.
3.2.1 Knowledge Concepts Extraction
Knowledge concepts used to be extracted by experts from domain texts. However, the rapid growth of the internet has caused conventional KCs extraction no longer meet the demand for handling large volumes of online data or updating domain knowledge in a timely fashion. To solve the problem, considerable researches aim to develop ML techniques to automate the KCs extraction. The existing ML methods for KCs extraction can be classified into two categories: supervised and unsupervised. The supervised extraction leverages classification models to categorize the text contents into numerous concepts (Firoozeh et al., 2020). However, this technique calls for annotated samples and its performance is heavily relied on the annotation qualities. On the other hand, unsupervised extraction recently dominated the recent KCs extraction because of the appearance of pre-trained language models. The initial unsupervised extraction was built with grammatical rules and semantic rules (Szwed, 2015; Stanković et al., 2016). For example, the Part-of-Speech (POS) tag rules such as Noun+Noun and [Adj or Noun]+Noun are commonly used to discern the KCs. Dictionary and Statistical based methods are later proposed to solve the low recalling issue of the rule-based method. Dictionary based methods (Gong et al., 2017; Levow et al., 2005) utilize a pre-defined concept dictionary to compare words or phrases in the text and extract relevant ones with similarity metrics. Statistical based methods (Boudin, 2018; Bougouin et al., 2013) analyze the word frequency and co-occurrence patterns, and use the rank of statistical features to identify KCs. The semantic-based extraction methods (Xiong et al., 2021; Huang et al., 2023) are the recent proposed ones which utilize the pre-trained models, e.g., word2vec (Mikolov et al., 2013) and BERT (Devlin et al., 2018), to obtain semantic representation of each candidate and apply post-processing filter models to extract target KCs. Because of the deeper meaning and contextual information captured by the pre-trained models, semantic-based extraction methods present the best result in both precision and recall among all unsupervised methods.
3.2.2 Pre-requisite Prediction
The pre-requisite prediction focuses on determining the relationship between two knowledge concepts. To be specific, for a concept pair $(A,B)$, there are four possible relations between them: (1) $A$ is a prerequisite of $B$; (2) $B$ is a prerequisite of $A$; (3) the two concepts are related, but they do not have any prerequisite relation; and (4) the two concepts are unrelated (Talukdar & Cohen, 2012). To simplify the problem, prior studies usually converted this task into a binary classification problem, which can be defined as:
$\mathrm{Preq}(B,A)=\begin{cases}1,A\mathrm{\ is\ a\ prerequisite\ of\ B}\\ 0,\mathrm{otherwise}\end{cases}$
The key components of pre-requisite prediction models are concept representation encoder and classification function. As concepts are usually associated with their description text, the text encoders are commonly employed as concept representation encoder. For the traditional methods, researches (Meng et al., 2016; Gasparetti, 2022) use bag-of-words with word frequencies or TF-IDF weights as concept representations. With the emergence of advanced text representation techniques, such as word2vec (Mikolov et al., 2013) and BERT (Devlin et al., 2018), recent approaches prefer to adopt these pre-trained embedding to capture the high-level semantic similarities between the concept pairs (Bai et al., 2021). The classification function is designed to map the representations of paired concepts to a scalar number which represents the likelihood of the pre-requisite relationship between two concepts. For simplicity, the inputs of classification function are usually the concept representations, but works (Bai et al., 2021) also explore to include manual features like hyperlink count, jaccard similarity and text categories. The common options for classification function include inner product, logistic regression and Multi-Layer Perceptron (MLP). To train pre-requisite prediction models, most studies use the classification-based losses such as binary cross entropy, but there are also studies (Liu et al., 2016) treating the problem as a ranking problem, which leverage the triplet loss.
3.3 Learning Path (LP) Generation
The individual differences cause the uniqueness in each learner’s optimal Learning Path (LP). With knowing learners’ characteristics, including preference, targets, and abilities, AL systems use the following ML methods to generate personalized learning paths, which are shown as Fig. 4.
Figure 4: Learning Path Generation Diagram.
3.3.1 Learner-based LP Generation
Learner-based LP focuses on creating a customized learning path by analyzing characteristics of the learning behavior observed during the learning process. Lin et al. (2013) first develop a personalized innovative learning system with decision tree. It considers students’ demographic information such as age, gender, major and creativity scores during LP recommendation process. Basu et al. (2013) and Salehi & Kamalabadi (2013) take more learner features, such as preference, academic performance, credit requirements, time availability into account, and leverage mixed, weighted, and cascade hybrid methods to form the final recommended learning paths. Salehi et al. (2014) propose a learning material recommendation system framework based on sequential pattern mining and multidimensional attribute-based Collaborative Filtering (CF). Yang & Dong (2017) propose a LP model that allows learning activities and the assessment criteria of their learning outcomes to be explicitly formulated by Bloom’s Taxonomy. Bendahmane et al. (2017) present a competence based approach derived from learning data, learner’s characteristics and their expectations. In this approach, learners were clustered and traced, and finally proper learning paths were presented. Dwivedi et al. (2018) introduce LPRS, which considers learners’ learning styles and knowledge levels and generates a learning path through a variable-length genetic algorithm.
3.3.2 Structure-based LP Generation
As knowledge concepts have strong logical sequences, many LP generation algorithms also exploit the relationship between knowledge elements to provide learners with an efficient learning path to their targets. Chu et al. (2011) propose a LP generation algorithm based on ontology. In this method, the ontology is established based on relationships between knowledge elements, and then a learning path is recommended according to the relationship between them in the ontology. Ouf et al. (2017) propose the framework for smart e-learning ecosystem using ontology and SWRL. Tam et al. (2014) present an explicit semantic analysis, followed by enhancing the ontology analysis through concept clustering, and applied an optimizer to find an optimal learning path of involved concepts or modules. Shmelev et al. (2015) combine the learner’s model with domain knowledge ontology of learning resources for matching, and generate personalized recommendation through the semantic recommendation algorithm. Al-Hassan et al. (2015) propose a hybrid semantic enhanced recommendation approach by combining the new Inferential Ontology-based Semantic Similarity (IOBSS) measure and standard item-based CF approach. Tseng et al. (2017) construct the concept map for adaptive learning and provide educational recommendations for individual students.
4 Bringing GenAI to Adaptive Learning

### v3

@@ Abstract
The recent surge in generative AI technologies, such as large language models and diffusion models, has boosted the development of AI applications in various domains, including science, finance, and education. Concurrently, adaptive learning, a concept that has gained substantial interest in the educational sphere, has proven its efficacy in enhancing students’ learning efficiency. In this position paper, we aim to shed light on the intersectional studies of these two methods, which combine generative AI with adaptive learning concepts. By presenting discussions about the benefits, challenges, and potentials in this field, we argue that this union will contribute significantly to the development of the next-stage learning format in education.
@@ 1 Introduction
Generative AI (GenAI) refers to artificial intelligence models that generate new content, including text, images, audio, and video, based on the patterns and information they have learned from the given training data (Cao et al., 2023). Unlike the traditional Machine Learning (ML) algorithms, which focus on analyzing and interpreting data, GenAI is designed to create new, original outputs and thus able to solve more challenging problems. With the recent success of GenAI technologies, such as large language models and diffusion models, in generating human-like outputs to a wide range of problems in Natural Language Processing (NLP) (Bubeck et al., 2023) and Computer Vision (CV) (Ramesh et al., 2022), the development of applying GenAI to real-world problems is rapidly increasing, which spreads from science (Walters & Murcko, 2020; Lopez et al., 2020), finance (Rane, 2023; Brühl, 2023) to education (Cooper, 2023; Baidoo-Anu & Owusu Ansah, 2023).
@@
Why This Position Paper? The overwhelming performance of GenAI over ML algorithms in challenging data analysis problems provides new opportunities to current ML-based AL systems. By leveraging the advanced capabilities of GenAI in creating human-level responses to problems like reasoning, the new AL systems will be able to finish more complicated AL tasks, such as dynamic learning path planning. Although some pioneering works have started to explore the usage of GenAI for education purposes (Dan et al., 2023; Xiao et al., 2024), there is no systematic review and discussion focusing on bringing GenAI to AL problems. As adaptive learning is increasingly recognized as a significant and promising direction for the future of education (Park & Lee, 2013), in this position paper, we propose to review the related studies and call attention to the opportunities in combining GenAI with AL. Furthermore, by conducting comprehensive discussions on the benefits, challenges, and threats, we hope to shed light on future research in this direction and push AL into the next chapter.
@@
Our Contributions. Our position in this paper is that bringing GenAI to AL will present both advantages and challenges to current ML-based AL systems, creating numerous opportunities and original topics for future educational development. To support this claim, we (1) conduct a comprehensive literature review on existing ML methods and GenAI works on adaptive learning, (2) explore the potential benefits and summarize ongoing industrial practices, (3) present in-depth discussions about the challenges and opportunities from the pedagogical perspective, and further extend these considerations to their broader educational impacts.
@@
Related Works. The impressive performance of GenAI, especially the Large Language Models (LLMs), in generating human-like responses to complicated requests (Bubeck et al., 2023) encourages the recent research in exploring GenAI for education purpose (Dan et al., 2023; Kieser et al., 2023). To provide researchers with a broad overview of the domain, numerous exploratory and survey papers have been proposed. For example, Qadir (2023); Rahman & Watanobe (2023) and Rahman & Watanobe (2023) conclude the applications of ChatGPT to engineering education by analyzing the responses of ChatGPT to the related pedagogical questions. Jeon & Lee (2023) and Mogavi et al. (2023) collect the opinions from different ChatGPT user groups, e.g., educator, learner, researcher, through in-person interviews, online post replies and user logs, and conclude the practical applications of LLMs in education scenarios. Baidoo-Anu & Owusu Ansah (2023) and Zhang & Tur (2023) focus on the literature review over the published papers and summarize the progress of the area with structured tables. Although the above works have covered a wide range of existing applications of GenAI in education scenario and provided their long-term visions for future studies, we argue that our position paper is still valuable as none of them focuses on the direction of bringing GenAI to AL. Compared to the general problems such as applying GenAI for education, we propose to present a more concentrated summarization and in-depth discussion toward the adaptive learning topic in this paper.
@@ 2.2 Generative AI (GenAI)
The origins of Generative AI can be traced back to the mid-1950s with the emergence of the concepts of artificial intelligence and machine learning. Unlike discriminative modeling, which has been the driving force of most progress in artificial intelligence, generative modeling problems are generally more difficult to tackle. Contributing to the rise of deep learning (LeCun et al., 2015), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Variational Auto-encoders(VAEs) (Kingma & Welling, 2013) first revolutionized image generation and became a cornerstone in the field of GenAI. Moreover, the recent emergence of Transformer (Vaswani et al., 2017), has marked another significant evolution in GenAI within the field of NLP, which has propelled GenAI into a new era. LLMs, such as OpenAI’s GPT series, are the most representative ones of these GenAI models. By taking advantage of the vast number of parameters, extensive pre-training text corpus, and advanced fine-tuning methods (Ouyang et al., 2022), LLMs have shown impressive capabilities in understanding context, generating coherent and contextually relevant responses (Bubeck et al., 2023). Apart from that, further studies (Brown et al., 2020) discover the strong few-shot (in-context) learning ability of LLMs, and such findings encourage the revolution in NLP solutions from task-specific methods to foundation models (Zhou et al., 2023). Despite the above progress, Generative Diffusion Models (GDMs) (Ho et al., 2020; Song et al., 2020) also bring improvements to the generation quality of images (Dhariwal & Nichol, 2021) and audio signals (Kong et al., 2020). Finally, by connecting GenAI models from different modalities, multi-modal GenAI models are proposed and have achieved promising results in directions like text-to-image (Ramesh et al., 2022), text-to-audio (Popov et al., 2021) and text-to-video generations (Wang et al., 2023a).
3 Bringing GenAI to AL
@@
In recent decades, ML plays a crucial role in AL by enabling the system to analyze learners’ behavior, build learning profiles, and provide personalized adaptions to learners. In following sections, we summarize changes brought by GenAI from two perspectives: (1) Empower existing algorithms, (2) Establishing novel directions. The relationships between GenAI and adaptive learning are summarized in Fig. 2.
Figure 2: An overview of bringing GenAI to AL.
3.1 Empower Existing Algorithms
Building students’ learning profile and then recommend suitable study materials are core functions for AL systems. In this section, we present a brief summarizing about existing works, and explain from different perspectives that GenAI will be an complementary helper to existing algorithms.
Figure 3: Empowering existing algorithms in AL system with GenAI.
3.1.1 Profile Building
The accurate profile information for different learner is an important foundation for any AL system. However, unlike profiles used in scenarios like commodity recommendation, profiles for adaptive learning focuses on tracking the learners’ knowledge proficiency, which cannot be observed directly. To solve this problem, prior works developed Cognitive Diagnosis Models (CDM) on the basis of cognitive science or psychometrics (Ackerman et al., 2003), and exploit different types of ML method, such as statistical learning models (Hambleton et al., 1991; De La Torre, 2009), neural networks (Cheng et al., 2019; Wang et al., 2020), to estimate the user’s knowledge proficiency based on the interaction history between the examinees and question. Commonly, CDM is applied in traditional paper-and-pencil tests, however, the one-for-all approach dismiss the difference between individuals and usually faces to the challenges in efficiency and accuracy. To deal with that, Computerized Adaptive Testing (CAT) is propose, and it aims at tailoring the selection of questions to each examinee’s level of proficiency, thereby maximizing the accuracy of the assessment while minimizing the test length (Liu et al., 2024). Due to the complexity of the optimization process, more advanced machine learning algorithms, such as reinforcement learning (Gilavert & Freire, 2022), active learning (Veldkamp & Verschoor, 2019; Bi et al., 2020), meta-learning (Ma et al., 2023a; Yu et al., 2023), are widely adopted by recent studies. Knowledge tracing (KT) is another popular research directions for learners’ profiles building, which also uses machine learning methods to build learners’ profiles. KT simplifies learner’s proficiency estimation as his/her next question correctness prediction (Abdelrahman et al., 2023), making it more flexible to diverse machine learning algorithms. The usual ML methods for KT include probabilistic graphical models (Corbett & Anderson, 1994), factor analysis models (Cen et al., 2006), and deep learning sequential models (Piech et al., 2015).
The success of above algorithms demonstrates the effectiveness of ML algorithms in adaptive learning, on the basis of that, we position that introducing GenAI could further exploit the potential of existing ML methods. First, the human-instructor-level performance of GenAI in composing questions in subjects like math (Lee et al., 2024), computer science (Prokudin et al., 2023) and language studies (Xiao et al., 2023) are widely recognized by recent studies. Following explorations by Adewumi et al. (2023) further verified the validness of using personalized question to enhance the engagement of student in test. These findings provide new opportunities for current CAT system in replacing pre-fixed question bank with dynamic generative pipelines, by which CAT can apply more flexible controls to questions and prior question bank exposure problem can be greatly resolved. Apart from that, leveraging GenAI to produce external knowledge for both KT and CAT data is also a valuable direction. As mentioned above, external knowledge like question knowledge concepts typologies is proven to be effective handlers in improving the performance of existing models. However, the external information requirement on data restricts the wide application of those new algorithms on different datasets. Contributing to powerful prior knowledge learnt by LLMs through the pre-training on large-scaled web corpus, using LLMs automatically add external knowledge, such as concept tags, to question becomes convenient, which will significantly encourage the emergency of stronger models on KT and CAT task. At last, by applying in-context learning (ICL) (Dong et al., 2022) and fine-tuning (Zhang et al., 2023a), LLMs can also directly be used as predictors. Pioneering studies by Neshaei et al. (2024) and Jung et al. (2024) observe that LLM based predictor has great potential in solving the ”cold-start” challenges for existing KT algorithms. Overall, the engagement of GenAI with existing profile building method brings new opportunities in the profile building researches of AL systems and the three above ways in leveraging LLMs for profile building is shown as Fig. 3.
3.1.2 Material Recommendation
Learning material recommendation is another important component of AL systems, which provides different user with the customized learning materials. Based on optimization objects and considering factors, current approaches can be classified into different categories: content based, collaborative filtering based, knowledge based, and hybrid solutions (Khanal et al., 2020). Content-Based (CB) methods use machine learning models to generate informative attributes for each learner (Shu et al., 2018). Based on these attributes, a rule-based system will exclude the contents and recommend the materials from remaining results (Kolekar et al., 2019). Collaborative Filtering (CF) make predictions about a learner’s interests based on preferences from many learners. In addition to preference, learner style (Bourkoukou et al., 2017) and skill levels (Han et al., 2016) are also considered by recent CF-based studies. Knowledge-based (KB) systems recommend items based on specific domain knowledge about how certain item features meet users’ needs and preferences (Aeiad & Meziane, 2019; Nitchot et al., 2019), and knowledge structures between learning contents are concerned by recent studies. At last, hybrid based solutions hybridizes the features of two or more above techniques to benefit from the strengths of each technique and to improve performance (Ghauth & Abdullah, 2010).
Although all above recommendation algorithms demonstrate good performance in their experiments, the application in real-world scenario still faces challenges, especially many of them require either delicate annotations or large scaled preference records (Khanal et al., 2020). In addition, as recent works prone to use deep learning models as their backbone model, many of these outputs are lack of intepretablity, which cannot provide valid support to learner’s inquiry on recommended materials. Fortunately, the recent progresses in using GenAI as knowledge enhancer (Gong et al., 2023) and predictor (Cui et al., 2022) for recommendation system provides good demonstrations in solve the issues above. To be specific, GenAI models like LLMs’ has powerful few-shot learning abilities, which enables it learn an effective predictor with limited annotated data (Brown et al., 2020). Besides that, LLMs’ strong prior knowledge can be leveraged as external resource to enhance the inputs for all textual factors considered within the recommendation process (Li et al., 2023). At last, the intermediate text before the outputs usually contain the reasoning process of LLMs, which automatically serve as good interpretation for the recommendation results (Chu et al., 2024). Based on these facts, we position that the introduction of GenAI will empower existing material recommendation algorithms with stronger adaptation capabilities and the recommendation process will be more transparent and convincing to the users. The usages of GenAI in material recommendation are detailed in Fig. 3.
3.2 Establishing Novel Directions
Apart from enhancing existing algorithms, GenAI’s generative characteristic and its emerged planning and imitation capability establish novel directions for AL studies. In this section, we focus on three representative ones and introduce changes brought by GenAI based method in each direction.
3.2.1 Content Creation
The high fidelity of generated results, such as image (Betker et al., 2023), video (Brooks et al., 2024), text (Achiam et al., 2023), demonstrates the great potential of GenAI in content creation. In addition, the emergent of techniques like instruction tunning (Zhang et al., 2023a) further facilitate the generating requirements by allowing users use verbal requirements to instruct LLMs during generation. Cooper (2023) explored the usages of GenAI models like GPT as instructor assistant to help create course outlines, compose a quiz focusing on some specific knowledge concepts, and even provide suggestion to teachers in organizing study plans for the specific lecture. Hu et al. (2024) leveraged GPT model to generate teaching plans in math for teachers and their evaluation results revealed the generated plans are achieving expert-level performance in multiple aspect, including problem chains organizations, teaching priority identifications and subject contents articulation. Besides textual learning contents, Pierce Burr & Young (2024) propose to use image-to-text generative model, DALLE-2 (Ramesh et al., 2022), to synthesis patient to educate doctors in training. By applying such approach, it will be possible to rapidly create photographic quality images without compromising patient confidentiality.
Figure 4: GenAI for Content Creation in AL.
As for adaptive learning, benefits brought by GenAI is not only limited to reduce the current workloads of instructors, but also provide new opportunities in creating learning contents aligning with learners’ interest and proficiency dynamically (Pesovski et al., 2024). Compared to static contents used by current AL systems, the dynamically created learning contents provide learner a more precised way in receiving knowledge. In addition, the real-time generation feature further empower AL systems to dynamically make adjustment based on changes happened in user’s knowledge proficiency during the learning process. At last, the recent progress in mutlimodal GenAI research provides good foundations to incorporate various modalities into GenAI models (Bewersdorff et al., 2024). In this case, the data formats of learning content will not be restricted by text only, the video and audio version contents can be helpful to improve the engagement of learner processes and provide convenient to people with special needs (Mallory, 2024). In a word, the emergence of GenAI is opening up new and creative ways to create contents for education and its usages for AL system are presented as Figure 4.
Figure 5: GenAI for Intelligent Agent in AL.
3.2.2 Intelligent Agent
Using GenAI model, LLMs, as intelligent agent is a popular direction in recent AI communities (Wang et al., 2024). The key idea of these researches is to leverage LLMs’ task planning potentials to solve complicated real-world problems through its initiative elaborations with external resources such as computer programs, search engine and other toolkit. On the basis of that, many advanced LLM-agent frameworks are proposed and achieves promising results in various type of problems such as math problem solving (Wu et al., 2024), program developing (Wu et al., 2023a) and question answering (Zhuang et al., 2024). In adaptive learning scenario, the direct application of intelligent agent is personal learning assistant chat-bot (Zarris & Sozos, 2023), which focuses on provide real-time support to student during their study. Contributing to the strong instruction following capability of GenAI models like LLMs, student can post his question and request directly through verbal text, and GenAI will be able to leverage all available resources to generate the answers aligning the with student’s needs. With incorporating techniques like prompt learning and fine-tuning, recent studies have successfully implemented learning chat-bots for essay writing (Han et al., 2024) and program learning (Smith et al., 2024). Apart from using it as tools only for student, agent-based framework can also be a powerful toolkit for instructors in adaptive learning. For example, providing personalized feedback on each student’s assignment is a burdensome work for teachers. Prior studies tried to use machine learning algorithms to automate the procedure but only focus on specific question types (Fagbohun et al., 2024). Contributing to the exceptional tool using ability, GenAI based algorithms is able to provide valid solutions to question with more general formats. Besides that, the exception planning capability of LLMs discovered by recent studies (Huang et al., 2024) also provides the vision that AL system can be organized and managed with the collaboration between different GenAI agents, and human instructors are free from the tedious routine works and focusing on providing supervisions on important decisions only.
3.2.3 Learning Simulation
Training models to produce indistinguishable samples with observed ones used to be a common training strategy for generative models like Generative Adversarial Networks (GAN) (Goodfellow et al., 2020). As the most advanced generative model, GenAI models have been proven be able to produce high-quality imitations of specific styled images (Zhang et al., 2023b) and text corpus (Tao et al., 2024). Moreover, contribute to the powerful in-context learning capabilities and conversational features of models like LLMs, using LLMs for role-playing and study the character of simulated roles has become an interesting topics in NLP communities (Shao et al., 2023). For adaptive learning, due to the privacy concerns and high collection costs, limited data size of student learning history is always the most common challenge faced by applying advanced machine learning models (Colchester et al., 2017). With leveraging the role-play ability of GenAI, it becomes possible to generate large scaled synthetic dataset with low cost. Inspired by recent studies in computer vision domain (Yeo et al., 2024), we are promising to see the performance gain added to current method with incorporating synthetic student data during the training. Furthermore, contribute to the strong profile following characteristics of the LLMs-based role-play algorithm (Shao et al., 2023), using it to test outcomes for different learning plans on specific student will be possible, which helps both instructor and AL systems to find optimal solutions from a more objective perspective.
Figure 6: GenAI for Learning Simulation in AL.
4 Impacts of GenAI on AL